OpenStudy (anonymous):

Write the expression as a single natural logarithm 3 In 3+3 In c

OpenStudy (anonymous):

first take 3 comn. then.. try

OpenStudy (anonymous):

ln M +ln N=?

OpenStudy (anonymous):

3(In 3+In c) =3In3c =In (3c)^3 answer In(27c^3)?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thanks